OpenStudy (bekkah323):

Money is invested in a savings account at 4.25% simple interest. After 1 year, there is $479.55 in the account. How much was originally invested?

OpenStudy (tkhunny):

What was the original deposit? If you do not know, then name it something.

OpenStudy (bekkah323):

i don't understand what your saying?

OpenStudy (tkhunny):

That's okay. It requires a little abstraction. P = The Amount of Money Originally Invested. Steve: How much money did you invest? Robert: P. Why so you ask? Once we have a name for it, and we can talk about it, it is much easier to build an equation. "Money is invested in a savings account at 4.25% simple interest. After 1 year, there is $479.55 in the account. How much was originally invested?" The standard formula for Simple Interest I = Prt We have r = 4.25% = 0.0425, t = 1, and hmmm...what is that $479.55? That is the total in the account after one year. It has two pieces. 1) The Original Deposit and 2) Interest Earned Add P to both sides of the equation. P + I = P + Prt Now we have P + I = 479.55. substitute all the known values. P + I = P + Prt Substituting r = 0.0425 P + I = P + P(0.0425)t Substituting t = 1 P + I = P + P(0.0425)(1) Substituting P + I = 479.55 479.55 = P + P(0.0425)(1) All we have left to do is solve for P. Can you do it?

OpenStudy (tkhunny):

Let's see what you got for P!

OpenStudy (bekkah323):

I follow you until the end equation, I don't know how to solve It?

OpenStudy (bekkah323):

@tkhunny

OpenStudy (tkhunny):

479.55 = P + P(0.0425)(1) Simplify 479.55 = P + P(0.0425) Distributive Property 479.55 = P(1 + 0.0425) = P(1.0425) Division Property 479.55 / (1.0425) = P How's that?

OpenStudy (tkhunny):

You do know that if you make me do all the work you have to work another problem for me so I can see if you are getting it!

OpenStudy (bekkah323):

Yes, thats fine, im studding for an exam so that would help alot

OpenStudy (bekkah323):

I Understand simplify and divide but on the distribute part I don't understand how you got there

OpenStudy (bekkah323):

@tkhunny

OpenStudy (tkhunny):

a(B+C) = aB + aC It also works the other way. aB + aC = a(B+C) That's all it is.